# Secure authentication via Quantum Physical Unclonable Functions: a review

### v1

@@ Abstract
Survey on QPUFs, QPUF, Authentication, Quantum memories
@@ 1 Introduction
Analogous to the classical case, token-based authentication schemes can be designed relying on Quantum PUFs (QPUFs) by exploiting the fingerprint set by the unpredictable outputs (responses) of the considered QPUF when being queried with new, i.e., unqueried, inputs (challenges). However, and distinctly, QPUFs, as formally defined in [18], accept quantum states as challenges, as well as they produce quantum states as responses. The first documented attempt of deriving what one might, at a glance, call a QPUF is found in [19]. Building on this work, [20] also proposes a PUF-based scheme involving quantum systems. We intentionally avoid calling the mentioned proposals QPUFs because, as it will become evident in Section 2, these fail to fall into the definition of a QPUF presented in [18]. A more adequate and accepted label for the schemes proposed in the aforementioned cited works is Quantum Readout PUF (QR-PUF). As we discuss in Section 4, these enjoy a reasonable sense of practicality, as opposed to QPUFs, but require additional ad-hoc assumptions and do not seem to take full advantage of the quantumness. That is, essentially, and as already pointed out in [21] for the specific work in [20], the quantum challenges and responses possible for QR-PUFs are typically mappable to classical information known by the certifier who, as opposed to that of QPUFs, becomes a trusted party.
@@
This article aims to provide the reader with a review on QPUFs, going in detail over all the aforementioned aspects within the complex development and evolution of an itself complex field of investigation. In Section 2.1 we revisit the theoretical framework introduced in [18], and we comment on some immediate implications of it when accepting certain assumptions commonly made for PUFs. In Section 2.2, we deliver a discussion on the current incompatibility between requirements and experimental possibilities for QPUF-based tokens, as well as the corresponding near-term perspectives. Later, in Section 3, we present and justify the article selection criteria used in this review, focusing on the most theoretically relevant QPUF proposals. In Section 4, we present our analysis of several articles, providing an underlying historical overview of QPUFs—illustrated as a QPUF timeline—, and a detailed discussion on information-theoretic approaches used for the study and characterization of QPUFs. Finally, Section 5 concludes the article.
@@ 2.1 Definitions and common assumptions
We begin by presenting the definition of a QPUF, originally introduced in [18]. It is important to note that this and the subsequent definitions have not been reproduced in an unaltered form. While our aim is not to change the meaning of the established concepts, we present slightly revised formulations that seek to enhance clarity and remain consistent with the original formalism.
@@
    $\delta_{u}$-uniqueness, ensuring that a sampled channel $\Lambda^{\delta_{r},\delta_{u},\delta_{c}}_{\{\lambda_{i}\}\text{-}\mathrm{QPUF}}$ is $\delta_{u}$-distinguishable (in the diamond norm [27, 28]) from any other sampled instance.
@@
On the almost-unitaricity requirement and the sampling problem:
In [18] (its Theorem 3) it is shown how the two last displayed requirements necessarily imply that $\Lambda^{\delta_{r},\delta_{u},\delta_{c}}_{\{\lambda_{i}\}\text{-}\mathrm{QPUF}}$ must be a unitary or a negligibly, with respect to certain security parameters, non-unitary channel. In parallel, the uniqueness requirement mandates the existence of a quantum-channel sampling procedure available to the verifier party, and establishing a proper fingerprint.
@@ 2.2 QPUF requirements and today’s possibilities
    Color centers: In materials like diamond, defects in the crystal lattice, such as vacancies or the presence of foreign atoms, can give rise to color centers. These defects form energy-level structures that are optically addressable and suitable for qubit implementation. Unlike trapped atoms, color centers are inherently confined within the solid-state lattice, eliminating the need for external trapping mechanisms. This simplifies the system design and allows for the integration of nanostructures around the qubit. However, the solid-state environment introduces decoherence, primarily due to interactions with other lattice defects and background noise.
@@
    Fiber loops: This approach represents a fundamentally different type of quantum memory, where the photon is stored by circulating it through a long optical fiber loop. The primary advantage is its simplicity, no quantum operations or auxiliary quantum systems are required to store the information. However, this method has notable limitations: photon loss accumulates over time as the photon travels through the fiber, and retrieval is not on-demand but strictly determined by the fixed length of the loop.
@@ 3 Methodology
Figure 1: Graph relation to [18] (left) and to [94] (right). We observe the collision of co-citations from the right graph into the left one (highlighted in red). The years of the co-related papers are specified.
In this review, we have selected the works of QR-QPUF [94], QPUF [18, 26], and Hybrid PUFs (HPUFs) [17] as our baseline articles, as they introduce novel theoretical perspectives related to the QPUF topic. It is important to note a selection bias in our choice of baselines, as we primarily focus on approaches that emphasize theoretical protocol innovation. Furthermore, for our analysis, we have set the acceptance threshold value to $l=0.1$. The outcome of the article selection is presented in Table  4 within the Appendix, where baseline articles are highlighted in yellow, and accepted articles for review are shaded in light gray. To better illustrate the relationships between articles, Figure  1 shows the co-citation network, with the article [18] on the left and [94] on the right, each serving as a baseline reference.
@@ 4.1 QR-PUFs
The first QR-PUF was coined in [19]. In this work, the author assumes the existence of a physically unclonable device capable of performing quantum unitary evolutions. Two further assumptions are, first, that distinct sampled unitaries are unique and distinguishable, and second, that attackers cannot emulate such unitary evolution with a sufficiently small time delay. Under these assumptions, two protocols are derived for remote authentication, each assuming different measurement capabilities at the verifier side. Moreover, another protocol achieving both unidirectional and mutual authentication for a Quantum Key Distribution (QKD) scheme is proposed.
@@ 4.2 QPUFs
    Firstly, the article [30] proposes using a QPUF device for client-server authentication. This work includes two authentication protocols with distinct hardware requirements: one being suitable for server authentication, with a Low Resources Verifier (LRV), and the other being suitable for client authentication, with a High Resources Verfier (HRV). The former protocol introduces major variations to the initial scheme found in [18], owning extra steps that include QPUF-device exchange and quantum state shuffling. At the verifier side, no quantum measurements are needed to be carried out, but quantum memories are still required. As for the latter defined protocol, the main novelty introduced is that it allows to choose between two different testing algorithms, one relying on the ordinary swap test, as originally conceived, and the other one relying on the so-called generalized swap (gswap) test.
@@ 4.3 Hybrid PUFs
In an attempt to gain practicality, the authors in [100] introduce the idea of combining classical and weak classical PUFs with a quantum encoding exploiting non-orthogonal states. Such construction, which they label as quantum lock, allows for having a Hybrid PUF (HPUF) that supports challenge reusability and is secure against attackers that have fully characterized the underlying classical PUF. On the other hand, both the offline and online protocols presented in [17] constitute two other HPUF models that, instead of relying on non-orthogonal state preparation, exploit the properties of maximally entangled states.
The above proposals represent a step towards requiring more realistic resources. For instance, they neither require quantum memories nor Haar randomness. Nevertheless, neither of the two articles discusses the effect of quantum noise or the threat posed by phishing-attack schemes, which can presumably compromise their security with the following strategy: the attacker impersonates the client first, in order to receive a challenge, and the server later, in order to obtain a valid response from the client and redirect it to the server.
@@ 4.4 Information-theoretic analysis of QPUFs
Information-theoretic methods also underpin secure key generation and data storage using QPUFs [105, 106]. The core insight is that QPUFs’ intrinsic randomness supplies high entropy that can be harvested into secret keys with provable secrecy. These high-entropy outputs can be processed (e.g., via privacy amplification) to yield secret keys. Information-theoretic analysis precisely quantifies how many secret bits can be extracted and how much information an adversary can learn. The secret key rate was shown to be related to the FAR, thereby establishing that a higher secret key rate extracted from a QPUF corresponds to a tighter bound on the adversarial success probability [104]. In [101, 107] secret key capacities extractable from QPUF outputs under privacy leakage have been bounded via mutual-information expressions. Extensions to secure data storage and message identification (detecting if a specific message is present instead of full decoding) via QPUFs were further characterized in [108] and [109], wherein identification capacity was shown to scale double-exponentially with the QPUF output length under idealized conditions. Importantly, these entropy-based bounds hold regardless of adversarial computational power. An information-theoretic argument guarantees that even a quantum-enabled adversary cannot reduce the QPUF’s inherent entropy below its limit. In turn, the QPUF’s intrinsic randomness is tied directly to provable secrecy: for example, any key derived from the QPUF can be made statistically independent of public helper data. These security guarantees rely on the concept of privacy leakage, which requires that public helper data reveal negligible information about the QPUF’s output. This requirement is especially critical in quantum settings, as compromise of QPUF output could lead to irreversible identity theft.
@@
All of these results rely on idealizations, i.i.d. noise, perfect challenge-response correlations, and unbounded blocklength, that abstract away practical limitations such as hardware imperfections, correlated errors, and finite‐length effects. The necessity of bridging the gap between theoretical information‐theoretic security proofs and realistic implementations has been emphasized in recent discussions [110, 111], and remains an important direction for future work. Despite these caveats, the information-theoretic framework provides a rigorous, implementation-independent foundation for the analysis of QPUF‐based cryptographic protocols, ensuring robustness even against adversaries endowed with quantum capabilities.
@@ 5 Conclusions
In this review article, we explore the concept of Quantum Physical Unclonable Functions (QPUFs) from both theoretical and practical perspectives. The initial theoretical framework serves as a crucial starting point for formalizing the essential criteria a QPUF must satisfy. However, while conceptually well defined, this model faces substantial experimental limitations that currently hinder its practical deployment. Among the most pressing challenges are the handling of quantum noise and the development of an effective and reliable QPUF sampler. However, a range of studies have emerged that build upon this foundational framework, either by strengthening its security properties or by relaxing some of its more demanding practical constraints. These contributions offer valuable insights into how theoretical models might evolve towards implementable forms.
@@ Acknowledgements
The authors gratefully acknowledge financial support from the Federal Ministry of Research, Technology and Space of Germany (BMFTR) within the program Souverän. Digital. Vernetzt. This work was supported by the joint project 6G-life, project identification numbers 16KISK002 and 16KISK263. Furthermore, VG and JN acknowledge additional support from BMFTR through projects 16KISQ039, 16KISQ077, and 16KISQ168, as well as from the German Research Foundation (DFG) under project NO 1129/2-1. CD and PJF received further funding from BMFTR projects 16KISQ169, 16KIS2196, 16KISQ038, 16KISR038, 16KISQ170, and 16KIS2234. HB and KN were supported by BMFTR projects 16KISQ037K, 16KISQ077, 16KISQ093, 16KISR026, 16KIS2195, and 16KIS1598K. TS, GP, and MB acknowledge support from the BMBF projects QPIS (No. 16KISQ032K) and DINOQUANT (No. 13N14921), as well as from the European Commission through the ERC Starting Grant project QUREP (No. 851810).

### v2

@@ Abstract
Review on QPUFs, QPUF, Authentication, Quantum memories
@@ 1 Introduction
Analogous to the classical case, token-based authentication schemes can be designed relying on Quantum PUFs (QPUFs) by exploiting the fingerprint set by the unpredictable outputs (responses) of the considered QPUF when being queried with new, i.e., unqueried, inputs (challenges). However, and distinctly, QPUFs, as formally defined in [18], accept quantum states as challenges, as well as they produce quantum states as responses. The first documented attempt of deriving what one might, at a glance, call a QPUF is found in [19]. Building on this work, [20] also proposes a PUF-based scheme involving quantum systems. We intentionally avoid calling the mentioned proposals QPUFs because, as it will become evident in Section 2, these fail to fall into the definition of a QPUF presented in [18]. An arguably more adequate label for the schemes proposed in the aforementioned cited works is Quantum Readout PUF (QR-PUF). As we discuss in Section 4, these enjoy a reasonable sense of practicality, as opposed to QPUFs, but require additional ad-hoc assumptions and do not seem to take full advantage of the quantumness. That is, essentially, and as already pointed out in [21] for the specific work in [20], the quantum challenges and responses possible for QR-PUFs are typically mappable to classical information known by the certifier who, as opposed to that of QPUFs, becomes a trusted party.
@@
This article aims to provide the reader with a review on QPUFs, going in detail over all the aforementioned aspects within the complex development and evolution of an itself complex field of investigation. In Section 2.1, we revisit the theoretical framework introduced in [18], and we comment on some immediate implications of it when accepting certain assumptions commonly made for PUFs. In Section 2.2, we deliver a discussion on the current incompatibility between requirements and experimental possibilities for QPUF-based tokens, as well as the corresponding near-term perspectives. Later, in Section 3, we present and justify the article selection criteria used in this review, focusing on the most theoretically relevant QPUF proposals. In Section 4, we present our analysis of several articles, providing an underlying historical overview of QPUFs—illustrated as a QPUF timeline—, and a detailed discussion on information-theoretic approaches used for the study and characterization of QPUFs. Section 5 provides an outlook on future research and Section 6 concludes the article.
@@ 2.1 Definitions and common assumptions
We begin by presenting the definition of a QPUF, originally introduced in [18]. It should be noted that this and the subsequent definitions are not reproduced verbatim. Rather, we introduce two deliberate adjustments. First, we allow for the possibility of incorporating more than one security parameter, thereby extending the scope of the formalism. Second, in order to keep the exposition clear and reader-oriented, we favor verbal formulations in place of fully technical mathematical expressions. These adjustments are not intended to alter the substance of the established concepts; instead, they are meant to offer a slightly different perspective that may facilitate the discussion while remaining faithful to the original framework.
@@
    $\delta_{u}$-uniqueness, ensuring that a generated channel $\Lambda^{\delta_{r},\delta_{u},\delta_{c}}_{\{\lambda_{i}\}\text{-}\mathrm{QPUF}}$ is $\delta_{u}$-distinguishable (in the diamond norm [27, 28]) from any other generated instance.
@@
On the almost-unitaricity requirement and the QPUF-generation problem:
In [18] (its Theorem 3) it is shown how the two last displayed requirements necessarily imply that $\Lambda^{\delta_{r},\delta_{u},\delta_{c}}_{\{\lambda_{i}\}\text{-}\mathrm{QPUF}}$ must be a unitary or a negligibly, with respect to certain security parameters, non-unitary channel. In parallel, the uniqueness requirement mandates the existence of a quantum-channel generating procedure available to the verifier party, and establishing a proper fingerprint.
@@ 2. Response as a token Holds the QPUF Stores a response (token) 1-shot verification
Remark 2.3.
On side-channel attacks:
While the information-theoretic security of QPUFs is robustly founded on quantum principles, their physical implementations introduce significant vulnerabilities to classical side-channel attacks that exploit unintentional information leakage such as power consumption, electromagnetic radiation, or photon statistics to infer the internal physical or quantum state evolution without directly measuring the quantum challenge–response. Mitigations include implementing randomized control sequences, cryogenic or electromagnetic shielding, and secure hardware design that minimizes state-dependent leakage. However, such countermeasures often introduce decoherence or impose high resource and calibration costs, thereby complicating the balance between physical security, stability, and scalability in realistic QPUF implementations.
@@ 2.2 QPUF requirements and today’s possibilities
An extended coherence time is crucial for the majority of quantum memory-based schemes. Nevertheless, there exist scenarios in which short-lived quantum tokens are sufficient, such as in digital signatures [45] and payment transactions [46], while other schemes do not require quantum memory[47, 48]. As summarized in Table 2, each physical platform exhibits a distinct combination of advantages and limitations. The realization of a quantum memory that meets the requirements of the QPUF-generated quantum token applications would require either a significant breakthrough within a single platform, such as the development of an engineered solid-state medium, or the implementation of a hybrid system that exploits the complementary strengths of different systems while suppressing their shortcomings.
@@
    Color centers: In materials like diamond, defects in the crystal lattice, such as vacancies or the presence of foreign atoms, can give rise to color centers. These defects form energy-level structures that are optically addressable and suitable for qubit implementation [68, 69, 70]. Unlike trapped atoms, color centers are inherently confined within the solid-state lattice, eliminating the need for external trapping mechanisms. This simplifies the system design and allows for the integration of nanostructures around the qubit. However, the solid-state environment introduces decoherence, primarily due to interactions with other lattice defects and background noise.
@@
    Fiber loops: This approach represents a fundamentally different type of quantum memory, where the photon is stored by circulating it through a long optical fiber loop. The primary advantage is its simplicity, no quantum operations or auxiliary quantum systems are required to store the information [72]. However, this method has notable limitations: photon loss accumulates over time as the photon travels through the fiber, and retrieval is not on-demand but strictly determined by the fixed length of the loop.
@@
Figure 1: In examining the graph relations of [18] to its co-citations, we observe overlapping co-citation patterns among several related nodes.
@@ 3 Methodology
In this review, we have selected the works of QR-QPUF [101], QPUF [18, 26], and Hybrid PUFs (HPUFs) [17] as our baseline articles, as they introduce novel theoretical perspectives related to the QPUF topic. It is important to note a selection bias in our choice of baselines, as we primarily focus on approaches that emphasize theoretical protocol innovation. Furthermore, for our analysis, we have set the acceptance threshold value to $l=0.1$. The outcome of the article selection is presented in Table  5 within the Appendix, where baseline articles are highlighted in yellow, and accepted articles for review are shaded in light gray. To better illustrate the relationships between articles, Figure  1 shows the co-citation network, with the article [18] serving as a baseline reference.
@@ 4.1 QR-PUFs
The first QR-PUF was coined in [19]. In this work, the author assumes the existence of a physically unclonable device capable of performing quantum unitary evolutions. Two further assumptions are, first, that distinct generated unitaries are unique and distinguishable, and second, that attackers cannot emulate such unitary evolution with a sufficiently small time delay. Under these assumptions, two protocols are derived for remote authentication, each assuming different measurement capabilities at the verifier side. Moreover, another protocol achieving both unidirectional and mutual authentication for a Quantum Key Distribution (QKD) scheme is proposed.
@@ 4.2 QPUFs
    Firstly, the article [30] proposes using a QPUF device for client-server authentication. This work includes two authentication protocols with distinct hardware requirements: one being suitable for server authentication, with a Low Resources Verifier (LRV), and the other being suitable for client authentication, with a High Resources Verifier (HRV). The former protocol introduces major variations to the initial scheme found in [18], owning extra steps that include QPUF-device exchange and quantum state shuffling. At the verifier side, no quantum measurements are needed to be carried out, but quantum memories are still required. As for the latter defined protocol, the main novelty introduced is that it allows one to choose between two different testing algorithms, one relying on the ordinary swap test, as originally conceived, and the other one relying on the so-called generalized swap (gswap) test.
@@ 4.3 Hybrid PUFs
In an effort to improve the practicality of weak PUFs, the authors in [107] propose the use of a quantum lock: a mechanism that combines classical and weak classical PUFs with a quantum encoding based on non-orthogonal states. The quantum lock enhances the security of weak PUFs by ensuring that even if an adversary has fully characterized the underlying classical PUF, the overall construction—referred to as a Hybrid PUF (HPUF)—remains secure. This approach also enables challenge reusability, a feature typically absent in PUF designs. On the other hand, both the offline and online protocols presented in [17] constitute two other HPUF models that, instead of relying on non-orthogonal state preparation, exploit the properties of maximally entangled states.
The above proposals represent a step towards requiring more realistic resources (see Table 4 for a key-feature comparison among all different PUFs existent today). For instance, they neither require quantum memories nor Haar randomness. Nevertheless, neither of the two articles discusses the effect of quantum noise or the threat posed by phishing-attack schemes, which can presumably compromise their security with the following strategy: the attacker impersonates the client first, in order to receive a challenge, and the server later, in order to obtain a valid response from the client and redirect it to the server.
Table 4: Comparison between different features of the different kinds of PUFs existent today.
CPUFs QR-PUFs QPUFs HPUFs
Trusted certifier Yes Yes No Yes
Security type Heuristic Heuristic Mathematically Heuristic
grounded
Commercially Yes No (to the best No (to the best of No (to the best
available today of our knowledge) of our knowledge) of our knowledge)
Experimentally Yes Yes No (to the best Yes
available today of our knowledge)
@@ 4.4 Information-theoretic analysis of QPUFs
Information-theoretic methods also underpin secure key generation and data storage using QPUFs [112, 113]. The core insight is that QPUFs’ intrinsic randomness supplies high entropy that can be harvested into secret keys with provable secrecy. Finally, the upper bound on the number of QPUFs that can be unambiguously identified within a network while maintaining reliable authentication and secure key generation is derived in [114]. These high-entropy outputs can be processed (e.g., via privacy amplification) to yield secret keys. Information-theoretic analysis precisely quantifies how many secret bits can be extracted and how much information an adversary can learn. The secret key rate was shown to be related to the FAR, thereby establishing that a higher secret key rate extracted from a QPUF corresponds to a tighter bound on the adversarial success probability [111]. In [108, 115] secret key capacities extractable from QPUF outputs under privacy leakage have been bounded via mutual-information expressions. Extensions to secure data storage and message identification (detecting if a specific message is present instead of full decoding) via QPUFs were further characterized in [116, 117, 118], wherein identification capacity was shown to scale double-exponentially with the QPUF output length under idealized conditions. Importantly, these entropy-based bounds hold regardless of adversarial computational power. An information-theoretic argument guarantees that even a quantum-enabled adversary cannot reduce the QPUF’s inherent entropy below its limit. In turn, the QPUF’s intrinsic randomness is tied directly to provable secrecy: for example, any key derived from the QPUF can be made statistically independent of public helper data. These security guarantees rely on the concept of privacy leakage, which requires that public helper data reveal negligible information about the QPUF’s output. This requirement is especially critical in quantum settings, as compromise of QPUF output could lead to irreversible identity theft.
@@
All of these results rely on idealizations, i.i.d. noise, perfect challenge-response correlations, and unbounded blocklength, that abstract away practical limitations such as hardware imperfections, correlated errors, and finite‐length effects. The necessity of bridging the gap between theoretical information‐theoretic security proofs and realistic implementations has been emphasized in recent discussions [119, 120, 121], and remains an important direction for future work. Despite these caveats, the information-theoretic framework provides a rigorous, implementation-independent foundation for the analysis of QPUF‐based cryptographic protocols, ensuring robustness even against adversaries endowed with quantum capabilities.
5 Outlook on future research
We conjecture that novel approaches may yet emerge within the QPUF literature. In analogy with the historical development of classical communication, one may envisage the possibility of a transformative idea in which quantum noise is harnessed as a resource rather than regarded exclusively as a hindrance [122, 123]. Since the field remains in an early stage of development and has not yet yielded fully satisfactory solutions to the problems it seeks to address, the exploration of new perspectives and methodologies appears both necessary and timely.
Future work is needed to define QPUF norms that can rigorously evaluate and certify QPUF devices for real-world deployment. Although significant progress has been achieved in the design and demonstration of QPUFs, the field currently lacks standardized frameworks for characterization, benchmarking, and interoperability across architectures such as photonic, superconducting, and ion-trap implementations. This absence impedes reproducibility, cross-platform comparison, and integration of QPUFs into broader quantum communication and cryptographic infrastructures. Existing evaluation metrics such as uniqueness, reliability, and unclonability remain inconsistently defined across experimental platforms, and no unified test protocols or certification criteria have been established. Future standardization efforts must therefore develop architecture-agnostic metrics, interoperable quantum–classical interfaces, rigorous verification methodologies, and seamless authentication across multi-vendor networks. Such standards will be essential for transitioning QPUFs from isolated laboratory prototypes to trusted, scalable components in quantum-secure hardware ecosystems. Furthermore, the high cost of quantum components currently restricts QPUFs to high-security domains like critical infrastructure and defense. Realizing QPUFs at scale will require coordinated advances across multiple disciplines: materials science and quantum engineering to enhance coherence, reduce noise, and integrate photonic or cryogenic components into compact, SWaP-C compliant packages.
@@ 6 Conclusions
In this review article, we explore the concept of Quantum Physical Unclonable Functions (QPUFs) from both theoretical and practical perspectives. The initial theoretical framework serves as a crucial starting point for formalizing the essential criteria a QPUF must satisfy. However, while conceptually well defined, this model faces substantial experimental limitations that currently hinder its practical deployment. Among the most pressing challenges are the handling of quantum noise and the development of an effective and reliable QPUF generator. However, a range of studies have emerged that build upon this foundational framework, either by strengthening its security properties or by relaxing some of its more demanding practical constraints. These contributions offer valuable insights into how theoretical models might evolve towards implementable forms.
@@
Before concluding this article, we want to further notice how, in this work, we have discussed how some quantum-based PUFs—such as QR-PUFs—can still be regarded as partly classical. Interestingly, the converse also holds: certain commercial implementations of classical PUFs, such as those by Crypto Quantique [124] and PUFsecurity [125], rely on fabrication-level variations originating from quantum phenomena, including tunneling in Complementary Metal–Oxide–Semiconductor (CMOS) or Metal–Oxide–Semiconductor Field-Effect Transistor (MOSFET) devices. Moreover, across all forms of PUFs, ensuring Truly Random Generation appears essential for security and unpredictability, as highlighted in recent industrial developments integrating quantum-based randomness [126, 127]. Quantum technologies can contribute significantly to this critical cryptographic primitive [128, 129].
@@ Acknowledgements
The authors thank Hadi Aghaee for valuable discussions. The authors further gratefully acknowledge financial support from the Federal Ministry of Research, Technology and Space of Germany (BMFTR) within the program Souverän. Digital. Vernetzt. This work was supported by the joint project 6G-life, project identification numbers 16KISK002 and 16KISK263. Furthermore, VG and JN acknowledge additional support from BMFTR through projects 16KISQ039, 16KISQ077, and 16KISQ168, as well as from the German Research Foundation (DFG) under project NO 1129/2-1. CD and PJF received further funding from BMFTR projects 16KISQ169, 16KIS2196, 16KISQ038, 16KISR038, 16KISQ170, and 16KIS2234. HB and KN were supported by BMFTR projects 16KISQ037K, 16KISQ077, 16KISQ093, 16KISR026, 16KIS2195, and 16KIS1598K. TS, GP, and MB acknowledge support from the BMBF projects QPIS (No. 16KISQ032K) and DINOQUANT (No. 13N14921), as well as from the European Commission through the ERC Starting Grant project QUREP (No. 851810).